# Computing the Intrinsic Delaunay Triangulation of a Closed Polyhedral Surface

### v1

@@ D.3 Interior of a thick biface
Proof D.19.
Let $t>1$. Assume $h_{S}(e)>12t\cdot(3+\log\lceil\ell(e)/s\rceil)$. We will prove $c_{S}(e)\geq t$, and this will prove the lemma. In $S$ there is a shortest path $p$ that intersects $e$ more than $12t\cdot(3+\log\lceil\ell(e)/s\rceil)$ times. Cut $e$ at its middle point. One of the two resulting sub-segments of $e$, say $f$, intersects $p$ more than $6t\cdot(3+\log\lceil\ell(e)/s\rceil)$ times. Partition $f$ into sub-segments $f_{0},f_{1},\dots,f_{n}$ with $n\leq 2+\log\lceil\ell(e)/s\rceil$, where the sub-segment $f_{0}$ contains the points $x\in f$ such that $\langle x\rangle_{e}\leq s/4$, and where for every $1\leq i\leq n$ the sub-segment $f_{i}$ contains the points $x\in f$ such that $2^{i-3}s\leq\langle x\rangle_{e}\leq 2^{i-2}s$. There is $0\leq i\leq n$ such that $p$ intersects $f_{i}$ more than $6t$ times. Then the relative interior of $p$ crosses $f_{i}$ twice (at least) in the same direction at points $x$ and $y$, such that the sub-segment of $f_{i}$ between $x$ and $y$ is shorter than $2^{i-4}s/t$, because $\ell(f_{i})\leq 2^{i-3}s$. Also $i\geq 1$ as no shortest path crosses $f_{0}$ twice, because $\ell(f_{0})<s/2$, and because the interior of $S$ is flat. In particular $\langle x\rangle_{e}\geq 2^{i-3}s$. Then $c_{S}(e)\geq t$ by Lemma D.16.
@@
Let us prove the claim. Consider a triangulation $T$ of $S$ with $n$ triangles, whose edges are all smaller than $L>0$. Cut each edge of $T$ into $2\lceil L/s\rceil$ equal-length segments, that we shall call doors. Each door is smaller than or equal to half the systole of $S$ so it is a shortest path. There are at most $6n\lceil L/s\rceil$ doors because $T$ has at most $3n$ edges. Each sub-segment $e_{1}$ of length $4L$ of $e_{0}$ contains in its relative interior three points $x_{0},x_{1},x_{2}$ in this order such that $x_{0}\notin p$, $x_{1}\in p$, and $x_{2}\notin p$ for some door $p$. The relative interior of $e_{0}$ intersects at least $30n\lceil L/s\rceil$ times doors this way, so there is a door $p$ intersected at least $5$ times by the relative interior of $e_{0}$. Then each intersection is a single point ($p$ and $e_{0}$ do not overlap). Two of those intersection points may be endpoints of $p$, but otherwise the relative interior of $p$ crosses the relative interior of $e_{0}$ at least three times. So $p$ crosses $e_{0}$ twice in the same direction, which proves the claim, and the proposition.
Proof D.22 (Proof of Proposition 5.3).
We have $h_{S}(e)=O(c_{S}(e)\cdot(1+\log\lceil\ell(e)/s\rceil))$ by Lemma D.18. Also $\ell(e)/s=O(c_{S}(e)\cdot n\cdot\lceil L/s\rceil^{2})$ by Lemma D.20. So $\log(\lceil\ell(e)/s\rceil)=O(1+\log c_{S}(e)+\log(n)+\log\lceil L/s\rceil)$. This proves the proposition.
@@
Proof E.1 (Proof of Lemma 5.14).
Assume that $e^{\prime}$ was not already an edge of $R_{A}^{1}$ for otherwise there is nothing to do. Then there is a good biface $B$ computed by step 2 of SimplifyTubes such that $e$ is one of the two interior edges of $B^{1}$. Also $B$ is thick, for $B$ has not been marked as inactive. So by Proposition D.7 there is a boundary edge $e$ of $B^{1}$ such that $c_{S}(e)\geq c_{S}(e^{\prime})-5$ and $\ell(e)\geq(1-4/c_{S}(e^{\prime}))\ell(e^{\prime})$.
@@
Proof F.1 (Proof of Lemma 5.16).
Otherwise one of the two loops of $X^{1}$ that constitute the boundary of $\mathcal{S}(X)$ is contractible in $S$. So this loop bounds a topological disk in $S$ by a result of Epstein [e-c2mi-66, Theorem 1.7]. The interior of the disk is flat, and its boundary is geodesic except possibly at the basepoint of the loop. This contradicts the formula of Gauss–Bonnet.
@@
We deduce Proposition 1.4 from Proposition 1.3, essentially by cutting out caps around the singularities in the interior of the surface, by applying Proposition 1.3 to the truncated surface, and by putting the caps back afterward. See Figure 6.
Figure 6: Cutting out a cap in the proof of Proposition 1.4.
Proof G.1 (Proof of Proposition 1.4).
@@
First construct in $O(n)$ time a triangular portalgon $T_{0}$ whose surface is $S$, as follows. Consider every singularity in the interior of $S$ (if any). This singularity is a vertex $v$ of $T^{1}$. Trace the boundary of the cap around $v$ in the faces of $T^{1}$. Then cut those faces along the trace, as in Figure 6. Afterward some polygons of $T_{0}$ may not be triangles, so cut each polygon of $T_{0}$ into triangles along shortcuts. Now remove the triangles corresponding to the caps from $T_{0}$, and let $T_{1}$ be the resulting triangular portalgon. The surface $\mathcal{S}(T_{1})$ is flat.
@@
The segment-happiness of $T^{\prime}$ and the happiness of $T^{\prime}$ do not differ by more than a constant factor because the polygons of $T^{\prime}$ are all triangles. Our second claim is that the segment-happiness of $T^{\prime}$, and thus the happiness of $T^{\prime}$, is bounded by $O(n\log(n)\log^{2}(r))$. To prove the second claim, we call cap path any shortest path in $S$ that lies in the closure of some cap. We call rogue path any shortest path in $S$ whose relative interior is disjoint from the closures of the caps. Every rogue path intersects every edge of $T^{\prime 1}$ at most $O(\log(n)\log^{2}(r))$ times, because the segment-happiness of $T_{1}$ is $O(\log(n)\log^{2}(r))$. Also every cap path intersects every edge of $T^{\prime 1}$ at most once. Now consider a shortest path $p$ in $S$. Then $p$ uniquely writes as a sequence $X$ of alternatively cap paths and rogue paths. Also, there cannot be two distinct cap paths $q_{0}$ and $q_{1}$ in $X$ that both lie in the same cap $C$. For otherwise any point of $q_{0}$ would be at distance at most $d/3$ from any point of $q_{1}$. Also the subpath of $p$ between $q_{0}$ and $q_{1}$ contains a rogue path that must leave the protected region around the central vertex of $C$, and is thus longer than d/2d/6=d/3d/2-d/6=d/3. That contradicts the fact that $p$ is a shortest path. We proved that there are at most $O(n)$ paths in $X$, each intersecting at most $O(\log(n)\log^{2}(r))$ times any given edge of $T^{\prime 1}$. This proves the second claim, and the proposition.
@@
Recall that we obtain our results within the real RAM model of computation described by Erickson, van der Hoog, and Miltzow [erickson2022smoothing]. It is an extension of the standard integer word RAM, with an additional memory array storing reals, and with additional instructions. The available instructions are described in [erickson2022smoothing, Table 1]. The arithmetic operations that can be performed by the machine on real numbers are addition, subtraction, multiplication, division, and square root.
@@
Figure 7: (Left) The polygons in the portalgon $T_{x}$. (Right) The polygons in the portalgon $D_{x}$, depending on whether $x=\lfloor x\rfloor$ or not.
To obtain our lower bound, we consider, for every $x\in(1,\infty)$, a particular representation of a portalgon $T_{x}$. See Figure 7. It has exactly two triangles. The first triangle has vertices $(0,0)$, $(1,0)$ and $(x,2)$. The second triangle is isometric to the first triangle, its vertices are $(1,0)$, $(x,2)$, and $(x+1,2)$. In $T_{x}$ the sides of the first triangle are matched with the corresponding sides of the second triangle. The surface $\mathcal{S}(T_{x})$ is a flat torus $S$, independent of $x$. We denote by $D_{x}$ the portalgon of the Delaunay tessellation of $S$.
@@
The proof of Lemma H.1 is deferred to Appendix H.1. A result similar to Lemma H.1 was proved by Blum, Shub, and Smale [blum1989theory, Section 4, Proposition 3] on a different machine, excluding the square root operation. We adapt their arguments to the machine described by Erickson, van der Hoog, and Miltzow [erickson2022smoothing], including the square root operation.
@@
Proof H.2 (Proof of Theorem 1.2).
The aspect ratio of $T_{x}$ is $O(x^{2})$ because the triangles of $T_{x}$ have maximum side length $O(x)$ and minimum height $\Omega(1/(1+x))$. We now prove that there is no real RAM program computing a representation of $D_{x}$ from $x\in(1,\infty)$ in $O((\log x)^{c})$ time for some $c\in(0,1)$. To do so we describe a real RAM program that, given $x$ and a representation of $D_{x}$, computes $\lfloor x\rfloor$ in constant time. This will prove the theorem by Lemma H.1. There are two cases. See Figure 7. Either $D_{x}$ has a unique polygon (in fact a rectangle), in which case $\lfloor x\rfloor=x$, so we return $x$. Otherwise $D$ has two triangles. In this case let $\Delta$ be any of these two triangles in our representation of $D_{x}$. Our triangle $\Delta$ has a unique side of length one. An orientation-preserving isometry of the plane displaces this unit length side to the segment between $(0,0)$ and $(1,0)$. The same isometry displaces the third vertex of $\Delta$ to a point $(u,v)$ such that $u=x-\lfloor x\rfloor$. We compute $u$ and return xux-u. This can be done using only the basic arithmetic operations of the real RAM on the coordinates of the vertices of $\Delta$. The theorem is proved.
@@
As already mentioned, a result similar to Lemma H.1 was proved by Blum, Shub, and Smale [blum1989theory, Section 4, Proposition 3] on a different machine, excluding the square root operation. We adapt their arguments to the machine described by Erickson, van der Hoog, and Miltzow [erickson2022smoothing], including the square root operation. This requires a few additional arguments.
@@
Note that without the square root operation, every function associated to a straight-line program would be rational, so it would flatten at at most one value. When including the square root operation however, such functions can flatten at several values: consider, for example, the function that maps each $x\in\mathbb{R}\setminus\{0\}$ to $\sqrt{x^{2}}/x$. This is why we need extra arguments compared to Blum, Shub, and Smale [blum1989theory, Section 4, Proposition 3].
Proof H.4 (Proof of Lemma H.3).
@@
By assumption $f$ is associated to a straight-line program $I_{1},\dots,I_{n}$. For each $i\in[n]$ we associate to the instruction $I_{i}$ some polynomial equation(s). For example each instruction of the form $x_{i}\leftarrow x_{j}\times x_{k}$ is associated to the single equation $x_{i}-x_{j}x_{k}=0$. And each instruction of the form $x_{i}\leftarrow\sqrt{x_{j}}$ is associated to the three equations $x_{i}^{2}-x_{j}=0$, $x_{i}>0$, and $x_{j}>0$. The other instructions are handled the same way. This defines at most $3n$ polynomial equations, each of degree at most two. The corresponding semi-algebraic subset $X\subset\mathbb{R}^{n+1}$ contains the possible values of the variables $(x_{0},\dots,x_{n})$ during an execution of the straight-line program. By definition the first co-coordinate projection maps $X$ to $D$ in a one-to-one manner, and this projection is continuous, so the number of connected components of $D$ is smaller than or equal to the number of connected components of $X$. And the latter is smaller than $(7n)^{3n}$ by a result of Milnor [milnor1964betti, Theorem 3].
Proof H.5 (Proof of Lemma H.1).
@@
Consider an integer $m\geq n+1$. By assumption, when executed with program $P$ on some input $x\in[n,m]$, the machine halts after at most $\alpha=(\log m)^{c}$ instructions. Which instructions are executed depends on $x$, and more precisely on the outcomes of the flow control instructions. Because at most $\alpha$ flow control instructions are executed, there are at most $2^{\alpha}$ possibilities for the outcomes of the flow control instructions. For each such possibility, the output (value of the first real register once the machine halts) is a function of the input (initial value of the first real register) that is associated to a straight-line program with at most $\alpha$ instructions. In particular, this function flattens at less than $(7\alpha)^{3\alpha}$ values by Lemma H.3, we will use this fact. Let $F$ contain all these functions, over all the possibilities for the outcomes of the flow control instructions. Now consider an integer $k\in[n,m-1]$. There is a non-empty open subset $O_{k}$ of the open interval $(k,k+1)$ such that the outcomes of the flow control instructions are the same for all inputs in $O_{k}$. So there is $f_{k}\in F$, whose domain contains $O_{k}$, such that $f_{k}(O_{k})=\{k\}$. In other words, $f_{k}$ flattens at $k$. Now recall that each $f_{k}$ flattens at less than $(7\alpha)^{3\alpha}$ values, so $(m-n)\leq 2^{\alpha}\cdot(7\alpha)^{3\alpha}$. As $m$ goes to $\infty$ this inequality becomes false, because $\alpha=(\log m)^{c}$ and $c<1$, a contradiction.

### v2

@@ D.3 Interior of a thick biface
Proof.
Let $t>1$. Assume $h_{S}(e)>12t\cdot(3+\log\lceil\ell(e)/s\rceil)$. We will prove $c_{S}(e)\geq t$, and this will prove the lemma. In $S$ there is a shortest path $p$ that intersects $e$ more than $12t\cdot(3+\log\lceil\ell(e)/s\rceil)$ times. Cut $e$ at its middle point. One of the two resulting sub-segments of $e$, say $f$, intersects $p$ more than $6t\cdot(3+\log\lceil\ell(e)/s\rceil)$ times. Partition $f$ into sub-segments $f_{0},f_{1},\dots,f_{n}$ with $n\leq 2+\log\lceil\ell(e)/s\rceil$, where the sub-segment $f_{0}$ contains the points $x\in f$ such that $\langle x\rangle_{e}\leq s/4$, and where for every $1\leq i\leq n$ the sub-segment $f_{i}$ contains the points $x\in f$ such that $2^{i-3}s\leq\langle x\rangle_{e}\leq 2^{i-2}s$. There is $0\leq i\leq n$ such that $p$ intersects $f_{i}$ more than $6t$ times. Then the relative interior of $p$ crosses $f_{i}$ twice (at least) in the same direction at points $x$ and $y$, such that the sub-segment of $f_{i}$ between $x$ and $y$ is shorter than $2^{i-4}s/t$, because $\ell(f_{i})\leq 2^{i-3}s$. Also $i\geq 1$ as no shortest path crosses $f_{0}$ twice, because $\ell(f_{0})<s/2$, and because the interior of $S$ is flat. In particular $\langle x\rangle_{e}\geq 2^{i-3}s$. Then $c_{S}(e)\geq t$ by Lemma D.7. ∎
@@
Let us prove the claim. Consider a triangulation $T$ of $S$ with $n$ triangles, whose edges are all smaller than $L>0$. Cut each edge of $T$ into $2\lceil L/s\rceil$ equal-length segments, that we shall call doors. Each door is smaller than or equal to half the systole of $S$ so it is a shortest path. There are at most $6n\lceil L/s\rceil$ doors because $T$ has at most $3n$ edges. Each sub-segment $e_{1}$ of length $4L$ of $e_{0}$ contains in its relative interior three points $x_{0},x_{1},x_{2}$ in this order such that $x_{0}\notin p$, $x_{1}\in p$, and $x_{2}\notin p$ for some door $p$. The relative interior of $e_{0}$ intersects at least $30n\lceil L/s\rceil$ times doors this way, so there is a door $p$ intersected at least $5$ times by the relative interior of $e_{0}$. Then each intersection is a single point ($p$ and $e_{0}$ do not overlap). Two of those intersection points may be endpoints of $p$, but otherwise the relative interior of $p$ crosses the relative interior of $e_{0}$ at least three times. So $p$ crosses $e_{0}$ twice in the same direction, which proves the claim, and the proposition. ∎
Proof of Proposition 5.2.
We have $h_{S}(e)=O(c_{S}(e)\cdot(1+\log\lceil\ell(e)/s\rceil))$ by Lemma D.8. Also $\ell(e)/s=O(c_{S}(e)\cdot n\cdot\lceil L/s\rceil^{2})$ by Lemma D.9. So $\log(\lceil\ell(e)/s\rceil)=O(1+\log c_{S}(e)+\log(n)+\log\lceil L/s\rceil)$. This proves the proposition. ∎
@@
Proof of Lemma 5.10.
Assume that $e^{\prime}$ was not already an edge of $R_{A}^{1}$ for otherwise there is nothing to do. Then there is a good biface $B$ computed by step 2 of SimplifyTubes such that $e$ is one of the two interior edges of $B^{1}$. Also $B$ is thick, for $B$ has not been marked as inactive. So by Proposition D.3 there is a boundary edge $e$ of $B^{1}$ such that $c_{S}(e)\geq c_{S}(e^{\prime})-5$ and $\ell(e)\geq(1-4/c_{S}(e^{\prime}))\ell(e^{\prime})$. ∎
@@
Proof of Lemma 5.11.
Otherwise one of the two loops of $X^{1}$ that constitute the boundary of $\mathcal{S}(X)$ is contractible in $S$. So this loop bounds a topological disk in $S$ by a result of Epstein [9, Theorem 1.7]. The interior of the disk is flat, and its boundary is geodesic except possibly at the basepoint of the loop. This contradicts the formula of Gauss–Bonnet. ∎
@@
We deduce Proposition 1.4 from Proposition 1.3, essentially by cutting out caps around the singularities in the interior of the surface, by applying Proposition 1.3 to the truncated surface, and by putting the caps back afterward. See Figure G.1.
Figure G.1: Cutting out a cap in the proof of Proposition 1.4.
Proof of Proposition 1.4.
@@
First construct in $O(n)$ time a triangular portalgon $T_{0}$ whose surface is $S$, as follows. Consider every singularity in the interior of $S$ (if any). This singularity is a vertex $v$ of $T^{1}$. Trace the boundary of the cap around $v$ in the faces of $T^{1}$. Then cut those faces along the trace, as in Figure G.1. Afterward some polygons of $T_{0}$ may not be triangles, so cut each polygon of $T_{0}$ into triangles along shortcuts. Now remove the triangles corresponding to the caps from $T_{0}$, and let $T_{1}$ be the resulting triangular portalgon. The surface $\mathcal{S}(T_{1})$ is flat.
@@
The segment-happiness of $T^{\prime}$ and the happiness of $T^{\prime}$ do not differ by more than a constant factor because the polygons of $T^{\prime}$ are all triangles. Our second claim is that the segment-happiness of $T^{\prime}$, and thus the happiness of $T^{\prime}$, is bounded by $O(n\log(n)\log^{2}(r))$. To prove the second claim, we call cap path any shortest path in $S$ that lies in the closure of some cap. We call rogue path any shortest path in $S$ whose relative interior is disjoint from the closures of the caps. Every rogue path intersects every edge of $T^{\prime 1}$ at most $O(\log(n)\log^{2}(r))$ times, because the segment-happiness of $T_{1}$ is $O(\log(n)\log^{2}(r))$. Also every cap path intersects every edge of $T^{\prime 1}$ at most once. Now consider a shortest path $p$ in $S$. Then $p$ uniquely writes as a sequence $X$ of alternatively cap paths and rogue paths. Also, there cannot be two distinct cap paths $q_{0}$ and $q_{1}$ in $X$ that both lie in the same cap $C$. For otherwise any point of $q_{0}$ would be at distance at most $d/3$ from any point of $q_{1}$. Also the subpath of $p$ between $q_{0}$ and $q_{1}$ contains a rogue path that must leave the protected region around the central vertex of $C$, and is thus longer than d/2d/6=d/3d/2-d/6=d/3. That contradicts the fact that $p$ is a shortest path. We proved that there are at most $O(n)$ paths in $X$, each intersecting at most $O(\log(n)\log^{2}(r))$ times any given edge of $T^{\prime 1}$. This proves the second claim, and the proposition. ∎
@@
Recall that we obtain our results within the real RAM model of computation described by Erickson, van der Hoog, and Miltzow [11]. It is an extension of the standard integer word RAM, with an additional memory array storing reals, and with additional instructions. The available instructions are described in [11, Table 1]. The arithmetic operations that can be performed by the machine on real numbers are addition, subtraction, multiplication, division, and square root.
@@
Figure H.1: (Left) The polygons in the portalgon $T_{x}$. (Right) The polygons in the portalgon $D_{x}$, depending on whether $x=\lfloor x\rfloor$ or not.
To obtain our lower bound, we consider, for every $x\in(1,\infty)$, a particular representation of a portalgon $T_{x}$. See Figure H.1. It has exactly two triangles. The first triangle has vertices $(0,0)$, $(1,0)$ and $(x,2)$. The second triangle is isometric to the first triangle, its vertices are $(1,0)$, $(x,2)$, and $(x+1,2)$. In $T_{x}$ the sides of the first triangle are matched with the corresponding sides of the second triangle. The surface $\mathcal{S}(T_{x})$ is a flat torus $S$, independent of $x$. We denote by $D_{x}$ the portalgon of the Delaunay tessellation of $S$.
@@
The proof of Lemma H.1 is deferred to Appendix H.1. A result similar to Lemma H.1 was proved by Blum, Shub, and Smale [3, Section 4, Proposition 3] on a different machine, excluding the square root operation. We adapt their arguments to the machine described by Erickson, van der Hoog, and Miltzow [11], including the square root operation.
@@
Proof of Theorem 1.2.
The aspect ratio of $T_{x}$ is $O(x^{2})$ because the triangles of $T_{x}$ have maximum side length $O(x)$ and minimum height $\Omega(1/(1+x))$. We now prove that there is no real RAM program computing a representation of $D_{x}$ from $x\in(1,\infty)$ in $O((\log x)^{c})$ time for some $c\in(0,1)$. To do so we describe a real RAM program that, given $x$ and a representation of $D_{x}$, computes $\lfloor x\rfloor$ in constant time. This will prove the theorem by Lemma H.1. There are two cases. See Figure H.1. Either $D_{x}$ has a unique polygon (in fact a rectangle), in which case $\lfloor x\rfloor=x$, so we return $x$. Otherwise $D$ has two triangles. In this case let $\Delta$ be any of these two triangles in our representation of $D_{x}$. Our triangle $\Delta$ has a unique side of length one. An orientation-preserving isometry of the plane displaces this unit length side to the segment between $(0,0)$ and $(1,0)$. The same isometry displaces the third vertex of $\Delta$ to a point $(u,v)$ such that $u=x-\lfloor x\rfloor$. We compute $u$ and return xux-u. This can be done using only the basic arithmetic operations of the real RAM on the coordinates of the vertices of $\Delta$. The theorem is proved. ∎
@@
As already mentioned, a result similar to Lemma H.1 was proved by Blum, Shub, and Smale [3, Section 4, Proposition 3] on a different machine, excluding the square root operation. We adapt their arguments to the machine described by Erickson, van der Hoog, and Miltzow [11], including the square root operation. This requires a few additional arguments.
@@
Note that without the square root operation, every function associated to a straight-line program would be rational, so it would flatten at at most one value. When including the square root operation however, such functions can flatten at several values: consider, for example, the function that maps each $x\in\mathbb{R}\setminus\{0\}$ to $\sqrt{x^{2}}/x$. This is why we need extra arguments compared to Blum, Shub, and Smale [3, Section 4, Proposition 3].
Proof of Lemma H.2.
@@
By assumption $f$ is associated to a straight-line program $I_{1},\dots,I_{n}$. For each $i\in[n]$ we associate to the instruction $I_{i}$ some polynomial equation(s). For example each instruction of the form $x_{i}\leftarrow x_{j}\times x_{k}$ is associated to the single equation $x_{i}-x_{j}x_{k}=0$. And each instruction of the form $x_{i}\leftarrow\sqrt{x_{j}}$ is associated to the three equations $x_{i}^{2}-x_{j}=0$, $x_{i}>0$, and $x_{j}>0$. The other instructions are handled the same way. This defines at most $3n$ polynomial equations, each of degree at most two. The corresponding semi-algebraic subset $X\subset\mathbb{R}^{n+1}$ contains the possible values of the variables $(x_{0},\dots,x_{n})$ during an execution of the straight-line program. By definition the first co-coordinate projection maps $X$ to $D$ in a one-to-one manner, and this projection is continuous, so the number of connected components of $D$ is smaller than or equal to the number of connected components of $X$. And the latter is smaller than $(7n)^{3n}$ by a result of Milnor [24, Theorem 3]. ∎
Proof of Lemma H.1.
@@
Consider an integer $m\geq n+1$. By assumption, when executed with program $P$ on some input $x\in[n,m]$, the machine halts after at most $\alpha=(\log m)^{c}$ instructions. Which instructions are executed depends on $x$, and more precisely on the outcomes of the flow control instructions. Because at most $\alpha$ flow control instructions are executed, there are at most $2^{\alpha}$ possibilities for the outcomes of the flow control instructions. For each such possibility, the output (value of the first real register once the machine halts) is a function of the input (initial value of the first real register) that is associated to a straight-line program with at most $\alpha$ instructions. In particular, this function flattens at less than $(7\alpha)^{3\alpha}$ values by Lemma H.2, we will use this fact. Let $F$ contain all these functions, over all the possibilities for the outcomes of the flow control instructions. Now consider an integer $k\in[n,m-1]$. There is a non-empty open subset $O_{k}$ of the open interval $(k,k+1)$ such that the outcomes of the flow control instructions are the same for all inputs in $O_{k}$. So there is $f_{k}\in F$, whose domain contains $O_{k}$, such that $f_{k}(O_{k})=\{k\}$. In other words, $f_{k}$ flattens at $k$. Now recall that each $f_{k}$ flattens at less than $(7\alpha)^{3\alpha}$ values, so $(m-n)\leq 2^{\alpha}\cdot(7\alpha)^{3\alpha}$. As $m$ goes to $\infty$ this inequality becomes false, because $\alpha=(\log m)^{c}$ and $c<1$, a contradiction. ∎